OpenStudy (anonymous):

u^2+2u-4=0 use the quadratic formula to solve.

OpenStudy (espex):

\[u=-(\sqrt{5}+1) u=(\sqrt{5}-1)\]

OpenStudy (anonymous):

\[u^{2}+2u-4=0\] \[u=-2 +/- \sqrt{2^{2}-4\left( 1 \right)\left( -4 \right)}/(2)\] \[=-2 +/-\sqrt{20}/(2)\] \[=-2+/-\sqrt{4 \times 5}/(2)\] \[=-2+/-2\sqrt{5}/(2)\] factor out a 2 from the top \[=2(-1+/-\sqrt{5})/(2)\] divide the top and bottom by 2 \[u=-1+/- \sqrt{5}\]

OpenStudy (anonymous):

from step 3 or anywhere else after that you can use a calculator to give you the answer in decimal form